OpenStudy (anonymous):

I need help please. Use the cofunction identities to find the angle theta that makes the statement true. sec(4theta+20degrees)=csc(2theta-7degrees)

OpenStudy (anonymous):

$$\sec x = \frac 1 {\cos x} \\ \csc x = \frac 1 {\sin x}$$

OpenStudy (anonymous):